# Compression Hacking: A Supplementary Perspective on Informatics Metric of Language Models from Geometric Distortion

### v1

@@ Abstract
Recently, the concept of “compression as intelligence” has provided a novel informatics metric perspective for language models (LMs), emphasizing that highly structured representations signify the intelligence level of LMs. However, from a geometric standpoint, the word representation space of highly compressed LMs tends to degenerate into a highly anisotropic state, which hinders the LM’s ability to comprehend instructions and directly impacts its performance. We found this compression-anisotropy synchronicity is essentially the “Compression Hacking” in LM representations, where noise-dominated directions tend to create the illusion of high compression rates by sacrificing spatial uniformity. Based on this, we propose three refined compression metrics by incorporating geometric distortion analysis and integrate them into a self-evaluation pipeline. The refined metrics exhibit strong alignment with the LM’s comprehensive capabilities, achieving Spearman correlation coefficients above 0.9, significantly outperforming both the original compression and other internal structure-based metrics. This confirms that compression hacking substantially enhances the informatics interpretation of LMs by incorporating geometric distortion of representations.
Compression Hacking: A Supplementary Perspective on Informatics Metric of Language Models from Geometric Distortion
Jianxiang Zang1, Meiling Ning2, Yongda Wei3, Shihan Dou1, Jiazheng Zhang1, Nijia Mo4, Binhong Li5, Tao Gui1thanks: Corresponding author., Qi Zhang1, Xuanjing Huang1∗ 1 Fudan University, 2 Beijing University of Posts and Telecommunications, 3 George Mason University 4 Shanghai University of International Business and Economics 5 Hong Kong University of Science and Technology (Guangzhou) zjxhgg@gmail.com, tgui@fudan.edu.cn
@@ 1 Introduction
  • The refined metrics exhibit significantly stronger alignment with LM’s comprehensive capabilities compared to original compression metric, thereby establishing a task-agnostic self-evaluation perspective for LMs. Anonymous codes available here.
@@ 2 Compression Hacking
In this section, we analyze the compression-anisotropy synchronicity in LM representations, where highly compressed LMs tend to exhibit word representations with strong anisotropy. Our investigation proceeds in two stages: First, we quantify both compression and anisotropy metrics by examining the internal structure of LM representations (covariance matrices). We then fit regression curves to model the relationship between anisotropy and compression, verifying it’s statistical significance. Second, through mechanistic analysis, we identify the underlying cause of this phenomenon, what we term “compression hacking”.
@@ 2.1 Preliminary: Differential Entropy based Compression Metric
Figure 2: Visualization of distribution of word representations and the eigenvalues across different models.
@@ 2.2 Anisotropy: The Geometric Property Correlated with Compression
The anisotropy of language models is a geometric property of representations that reflects the non-uniform distribution of semantics across different directions in the representation space Ethayarajh (2019); Cai et al. (2019); Demeter et al. (2020). Highly anisotropic representations hinder LMs’ ability to comprehend instructions Yu et al. (2022); Rudman and Eickhoff (2024), directly impairing their overall capabilities. We performed principal component analysis to visualize the word representation spaces of the aforementioned four models. As shown in Figure 2, we made the intriguing observation that models with higher compression levels consistently exhibited greater unevenness in their dimensional distributions, namely, higher anisotropy. This suggests a potential synergistic relationship between compression and anisotropy. If we can quantify this relationship and confirm its statistical significance, it could provide valuable guidance for refining compression metrics.
@@ 3.2 Evaluation Pipeline
In this section, we integrate the three refined metrics into a unified evaluation framework, which is a task-agnostic pipeline operating purely from a representational perspective. Our evaluation paradigm associates the sampled data batch $\mathcal{B}$ with a decision score $s=\mathcal{F}(\mathcal{B},f_{\text{LM}})$. The decision function $\mathcal{F}(\cdot)$ operates through two sequential processes: (1) the projection step extracts hidden representations $\mathbf{Z}^{(\bm{p})}=\mathcal{F}_{\text{Projection}}(\bm{p},f_{\text{LM}})$ for each data sample $\bm{p}\in\mathcal{B}$; (2) the decision step computes the batch-level score $s=\mathbb{E}_{\bm{p}\sim\mathcal{B}}\text{Metric}(\mathbf{Z}^{\bm{p}})$ based on the refined metrics. Notably, our dataset requirement specifies that the sample’s word representation space should effectively estimate the model’s complete word representation space given sufficient sampling, ensuring convergence of our proposed metrics. We discuss the impact of sampling size on metric convergence in Section D.
@@ A.1 Evaluation of Language Models
Recently, significant efforts have been devoted to exploring the mechanisms by which language models (LMs) process information internally, driving the development of LM self-evaluation Wei et al. (2024); Wang et al. (2024a, b) independent of specific tasks and model outputs. The concept of “compression as intelligence” has provided an information-theoretic internal evaluation perspective for language models, highlighting that the acquisition of world knowledge by language models is a denoising process Sutskever (2023); Deletang et al. (2023); Wei et al. (2024); Chen et al. (2025). Differential entropy of representations, as a classical information-theoretic measure, effectively quantifies the internal uncertainty of language models Chen et al. (2023a); Zhouyin and Liu (2023). Semantic volume Li et al. (2025) leverages representation-level differential entropy-aware compression metrics to offer a novel perspective for language model evaluation. However, related work has found that such compression can only model the scale of language models and fails to align with their capabilities Wei et al. (2024); Li et al. (2025).
@@ A.2 Anisotropy of Language Models
Anisotropy Razors To mitigate anisotropy in language models, existing research has proposed various solutions. Contrastive learning has emerged as a powerful tool for obtaining effective sentence representations, effectively reducing anisotropy by increasing the spatial distance between positive and negative samples Gao et al. (2021); Zhang et al. (2022b); Jiang et al. (2022). In this work, we employ post-processing methods applied directly to the representation space as baseline approaches for the anisotropy razor:
@@ Appendix D Implementation Details of the Evaluation Pipeline
For the projection dataset, we primarily collected 1,000 data samples from the pretraining corpus (Wiki Foundation (2025)) and the instruction-tuning dataset (Dolly-15k Conover et al. (2023)) to derive projection data. By sampling the word representations of these data points, we aim to estimate the full model’s representation space, ensuring the convergence of our metrics. Our pipeline defaults to sampling 800 data samples. Figure 8 illustrates the cumulative expected values of different metrics as the number of samples increases. We observe that all metrics converge relatively early to stable values, demonstrating that our refined metrics enable robust evaluation based on the provided projection dataset.

### v2

# Compression Hacking: A Supplementary Perspective on Informatics Properties of Language Models from Geometric Distortion

@@ Abstract
Recently, the concept of “compression as intelligence” has provided a novel informatics metric perspective for language models (LMs), emphasizing that highly structured representations signify the intelligence level of LMs. However, from a geometric standpoint, the representation space of highly compressed LMs tends to degenerate into a highly anisotropic state, which hinders the LM’s ability to comprehend instructions and directly impacts its performance. We found this compression-anisotropy synchronicity is essentially the “Compression Hacking” in LM representations, where noise-dominated directions tend to create the illusion of high compression rates by sacrificing spatial uniformity. Based on this, we propose three refined compression metrics by incorporating geometric distortion analysis and integrate them into a self-evaluation pipeline. The refined metrics exhibit strong alignment with the LM’s comprehensive capabilities, achieving Spearman correlation coefficients above 0.9, significantly outperforming both the original compression and other internal structure-based metrics. This confirms that compression hacking substantially enhances the informatics interpretation of LMs by incorporating geometric distortion of representations.
Compression Hacking: A Supplementary Perspective on Informatics Properties of Language Models from Geometric Distortion
Jianxiang Zang1, Meiling Ning2, Yongda Wei3, Shihan Dou1, Jiazheng Zhang1, Nijia Mo4, Binhong Li5, Tao Gui1thanks: Corresponding author., Qi Zhang1, Xuanjing Huang1∗ 1 Fudan University, 2 Beijing University of Posts and Telecommunications, 3 George Mason University 4 Shanghai University of International Business and Economics 5 Hong Kong University of Science and Technology (Guangzhou) jxzang25@m.fudan.edu.cn, tgui@fudan.edu.cn
@@ 1 Introduction
  • The refined metrics exhibit significantly stronger alignment with LM’s comprehensive capabilities compared to original compression metric, thereby establishing a task-agnostic self-evaluation perspective for LMs. Codes available here.
@@ 2 Compression Hacking
In this section, we analyze the compression-anisotropy synchronicity in LM representations, where highly compressed LMs tend to exhibitin-context representations with strong anisotropy. Our investigation proceeds in two stages: First, we quantify both compression and anisotropy metrics by examining the internal structure of LM representations (covariance matrices). We then fit regression curves to model the relationship between anisotropy and compression, verifying it’s statistical significance. Second, through mechanistic analysis, we identify the underlying cause of this phenomenon, what we term “compression hacking”.
@@ 2.1 Preliminary: Differential Entropy based Compression Metric
Figure 2: Visualization of distribution ofin-context representations and the eigenvalues across different models.
@@ 2.2 Anisotropy: The Geometric Property Correlated with Compression
The anisotropy of language models is a geometric property of representations that reflects the non-uniform distribution of semantics across different directions in the representation space Ethayarajh (2019); Cai et al. (2019); Demeter et al. (2020). Highly anisotropic representations hinder LMs’ ability to comprehend instructions Yu et al. (2022); Rudman and Eickhoff (2024), directly impairing their overall capabilities. We performed principal component analysis to visualize thein-context representation spaces of the aforementioned four models. As shown in Figure 2, we made the intriguing observation that models with higher compression levels consistently exhibited greater unevenness in their dimensional distributions, namely, higher anisotropy. This suggests a potential synergistic relationship between compression and anisotropy. If we can quantify this relationship and confirm its statistical significance, it could provide valuable guidance for refining compression metrics.
@@ 3.2 Evaluation Pipeline
In this section, we integrate the three refined metrics into a unified evaluation framework, which is a task-agnostic pipeline operating purely from a representational perspective. Our evaluation paradigm associates the sampled data batch $\mathcal{B}$ with a decision score $s=\mathcal{F}(\mathcal{B},f_{\text{LM}})$. The decision function $\mathcal{F}(\cdot)$ operates through two sequential processes: (1) the projection step extracts hidden representations $\mathbf{Z}^{(\bm{p})}=\mathcal{F}_{\text{Projection}}(\bm{p},f_{\text{LM}})$ for each data sample $\bm{p}\in\mathcal{B}$; (2) the decision step computes the batch-level score $s=\mathbb{E}_{\bm{p}\sim\mathcal{B}}\text{Metric}(\mathbf{Z}^{\bm{p}})$ based on the refined metrics. Notably, our dataset requirement specifies that the sample’sin-context representation space should effectively estimate the model’s completein-context representation space given sufficient sampling, ensuring convergence of our proposed metrics. We discuss the impact of sampling size on metric convergence in Section D.
@@ A.1 Evaluation of Language Models
Recently, significant efforts have been devoted to exploring the mechanisms by which language models (LMs) process information internally, driving the development of LM self-evaluation Wei et al. (2024); Wang et al. (2024a, b); Zang and Liu (2024a); Dou et al. (2024) independent of specific tasks and model outputs. The concept of “compression as intelligence” has provided an information-theoretic internal evaluation perspective for language models, highlighting that the acquisition of world knowledge by language models is a denoising process Sutskever (2023); Deletang et al. (2023); Wei et al. (2024); Chen et al. (2025). Differential entropy of representations, as a classical information-theoretic measure, effectively quantifies the internal uncertainty of language models Chen et al. (2023a); Zhouyin and Liu (2023). Semantic volume Li et al. (2025) leverages representation-level differential entropy-aware compression metrics to offer a novel perspective for language model evaluation. However, related work has found that such compression can only model the scale of language models and fails to align with their capabilities Zang and Liu (2023b); Wei et al. (2024); Li et al. (2025).
@@ A.2 Anisotropy of Language Models
Anisotropy Razors To mitigate anisotropy in language models, existing research has proposed various solutions. Contrastive learning has emerged as a powerful tool for obtaining effective sentence representations, effectively reducing anisotropy by increasing the spatial distance between positive and negative samples Gao et al. (2021); Zhang et al. (2022b); Jiang et al. (2022); Zang and Liu (2023a, 2024b). In this work, we employ post-processing methods applied directly to the representation space as baseline approaches for the anisotropy razor:
@@ Appendix D Implementation Details of the Evaluation Pipeline
For the projection dataset, we primarily collected 1,000 data samples from the pretraining corpus (Wiki Foundation (2025)) and the instruction-tuning dataset (Dolly-15k Conover et al. (2023)) to derive projection data. By sampling thein-context representations of these data points, we aim to estimate the full model’s representation space, ensuring the convergence of our metrics. Our pipeline defaults to sampling 800 data samples. Figure 8 illustrates the cumulative expected values of different metrics as the number of samples increases. We observe that all metrics converge relatively early to stable values, demonstrating that our refined metrics enable robust evaluation based on the provided projection dataset.